# Hypervelocity star observations constrain the Galactic Centre

### v1

@@ Abstract
Hypervelocity stars (HVSs) are stars which have been ejected from the Galactic Centre (GC) at velocities of up to a few thousand $\text{km}\,\text{s}^{-1}$. They are tracers of the Galactic potential and can be used to infer properties of the GC, such as the initial-mass function and assembly history. HVSs are rare, however, with only about a dozen promising candidates discovered so far. In this work we use a novel, highly efficient method to identify new HVS candidates in Gaia. This method uses the nearly radial trajectories of HVSs to infer their distances and velocities based on their position and Gaia proper motion alone. Through comparison of inferred distances with Gaia parallaxes and photometry we identified 600 HVS candidates with G<20 including the previously discovered S5-HVS1, out of which we obtained ground-based follow-up observations for 196 stars. As we found no new HVSs based on their radial velocity, we used detailed HVS ejection simulations to significantly improve previous HVS ejection rate constraints. In particular, the ejection rate of HVSs more massive than 1 $\mathrm{M_{\odot}}$ cannot be higher than $10^{-5}$ yr-1 at $2\sigma$ significance. Additionally, we predict that there are 5–45 unbound HVSs in the complete Gaia catalogue ($1\sigma$ interval), most of which will be main-sequence stars of a few M at heliocentric distances of tens to hundreds of kpc. By comparing our results to literature HVS candidates, we find an indication of either a time-dependent ejection rate of HVSs or a non-GC origin of many previously identified HVS candidates.
@@
pubyear: 2024pagerange: Hypervelocity star observations constrain the Galactic CentreD
@@ 2 HVS candidate selection
The main challenge when searching for HVSs is that they are extremely rare; there are for instance no HVSs with velocity greater than 700 km s-1 in the 34M sources in Gaia DR3 with radial velocities, that could be robustly identified (Marchetti et al., 2022). HVS candidates selected from observational catalogues for follow-up observations are therefore typically strongly contaminated by non-HVSs. Here we present our new, method to identify HVS candidates originating from the GC. In Section 2.1 we present the general principle we use to find HVSs, and in Section 2.2 we provide the precise selections on the Gaia catalogue for our observational campaign.
@@ 2.1 Radial trajectory
Let $\vec{R}$ be the vector from the GC to the star, and $\vec{V}$ the star’s velocity vector relative to the GC. We can write these down in terms of the observables
@@
Figure 1: A diagram of the position ($\vec{R}$) and velocity ($\vec{V}$), in blue, of an HVS relative to the GC. This can be expressed in terms of a combination of the star’s observed position and proper motion (pink) and measurements of the Sun’s position and velocity (black) along with the distance to the star ($D$) and its radial velocity ($V_{r}$).
@@
We convert $D_{\rm I}$ to implied parallax ($\varpi_{\rm I}$) to allow for convenient comparison to measurements by taking the inverse of $D_{\rm I}$. We calculate the uncertainty on $\varpi_{\rm I}$ assuming that the proper motion measurement is the only source of uncertainty. We make the simplifying assumption that the proper motion uncertainties are uncorrelated and since $\varpi\propto\mu$, we can approximate the uncertainty on $\varpi$ as
$\sigma_{\varpi}=\frac{\sqrt{\left[\vec{R}_{0}\cdot\left(\hat{n}\times\hat{\mu}_{\delta}\right)\sigma_{\mu_{\delta}}\right]^{2}+\left[\vec{R_{0}}\cdot\left(\hat{n}\times\hat{\mu}_{\alpha*}\right)\sigma_{\mu_{\alpha*}}\right]^{2}}}{\hat{n}\cdot\left(\vec{R}_{0}\times\hat{V}_{0}\right)},$ (7)
with $\hat{\mu}{{}_{\delta}}$ and $\hat{\mu}_{\alpha*}$ the basis vectors for the proper motion in declination (Dec) and right ascension (RA) respectively, and $\sigma_{\mu_{\delta}}$ and $\sigma_{\mu_{\alpha*}}$ the uncertainties on the proper motions in Dec and RA respectively (where we take $\sigma_{\mu_{\alpha}*}=\sigma_{\mu_{\alpha}}\cos^{2}\delta$). For a more accurate uncertainty analysis one would sample over the uncertainty on all observables.
@@ 2.2.1 Astrometric and kinematic selections
We start by using the astrometric and kinematic properties of stars to select potential HVSs. HVSs need to be very fast to prevent significant deviation from their otherwise radial trajectories by the torque on the orbit generated by the non-sphericity of the Galactic potential. For this reason, we only consider stars with total implied velocity, $V_{\text{I}}$, in the range $[800,3500]$ $\text{km}\,\text{s}^{-1}$. Secondly, we compare the implied parallax to the parallax measured by Gaia. We only select sources with implied parallax consistent with measured parallax within $2\sigma$, considering both the uncertainty on the measured and the implied parallaxes. We do this comparison to the parallax and not the distance since the parallax has Gaussian uncertainties.
@@ 2.2.3 Sky coordinate selections
We know that the distribution of HVSs on the sky should be anisotropic, even for an isotropic ejection mechanism in the GC due to on-sky projection. Moreover, we expect that the contamination of our HVS candidate sample is highest near the GC and anti-centre, where our selections are relatively ineffective. In addition, in the Galactic plane we expect the contamination to be higher than at high Galactic latitudes due to the high relative number of field stars to expected HVSs. In order to limit contamination by field stars we define the expected number of HVSs to number of candidates for different directions in the sky. For this purpose we use HEALPix [^2] http://healpix.sf.net/ ($nside=3$; Górski et al., 2005; Zonca et al., 2019) to divide the sky into equal area pixels. We determine the expected number of HVSs using the simulations described in Section 4. We only observe sources in HEALPix pixels where the ratio of the normalised number of expected HVSs to candidates is larger than five [^3] The HEALPix pixels we use are 0-19, 22-31, 35-42, 46-55, 58-62, 65, 66, 69-73, 78, 82-85, 93-96, 101, 102, 103 and 107 in the ring scheme based on RA and Dec..
@@
The selection steps up to this point leave us with only 600 HVS candidates with ${\rm G}<20$, the previously identified S5-HVS1 being one of the top candidates (Koposov et al., 2020). We provide an overview of the selections in Table 1.
@@ 3.4 Observed HVS candidates
Figure 3: Sky distribution of our HVS candidate sample. Black points indicate the stars with follow-up observations and the orange points the ones with either SDSS or LAMOST radial velocity measurements. For the grey stars we do not have radial velocity measurements.
@@
In Fig 4 we plot the implied radial velocity against the measured radial velocity.
@@ 4.1 Hills mechanism
where we neglected a multiplicative constant factor of order of unity which depends on the geometry of the encounter (see, e.g., fig.10 in Sari et al. (2009). In eq. 8, $m_{\text{c}}$ is the mass of the captured companion, $M_{\text{MBH}}$ in the case of the Galaxy is the mass of Sgr A ($4\times 10^{6}$ M, Eisenhauer et al., 2005; Ghez et al., 2008; Schödel et al., 2009; Gillessen et al., 2009; Boehle et al., 2016; Gillessen et al., 2017; Do et al., 2019; GRAVITY Collaboration et al., 2019), and $M$ the total mass of the progenitor binary (Sari et al., 2009; Kobayashi et al., 2012; Rossi et al., 2014). Most of the stars ejected through this mechanism are still bound to the Galaxy, but the ejections can occur at velocities in excess of 1000 $\text{km}\,\text{s}^{-1}$, putting stars on orbits unbound to the Galaxy.
@@ 4.2 Simulations
We start by creating an HVS progenitor population of binaries. These binaries are characterised by the primary zero-age main sequence mass ($m_{\text{p}}$), mass ratio ($q$), and the orbital semi-major axis ($a$) at the moment of disruption. The mass of the primary is drawn from a mass function (MF) which follows a power law distribution of the form $f(m_{\text{p}})\propto m_{\text{p}}^{-\kappa}$ in the range $[0.1,100]$ M. The secondary is then assigned a mass according to a mass ratio distribution of the form $f(q)\propto q^{\gamma}$, with $0.1\leq q\leq 1$. Lastly, the orbital semi-major axis are assigned assuming a binary orbital period distribution of the form $f(\log P)\propto(\log P)^{\pi}$. Different values for the three parameters $\kappa$, $\gamma$, and $\pi$ are explored with uniform priors to account for uncertainty in the HVS progenitor population. For $\kappa$ we use the interval $[0.3,3.3]$ to allow for both very top-heavy and bottom-heavy MFs. For $\gamma$ we use the interval $[-2,2]$ and for $\pi$ the interval $[0,2]$. We allow for the wide range in MF index because of the uncertainty on the IMF near the GC, alluded to in the introduction. IMF indexes of 0.45, 1.7, and a canonical 2.3 have all been claimed in Sgr A’s sphere of influence (Bartko et al., 2010; Lu et al., 2013; Löckmann et al., 2010, respectively). The parameter range on $\pi$ allows for a log-uniform period distribution (also known as Öpik’s law), or distributions with increasing fractions of wide binaries. The parameter range on $q$ encapsulates the results from different studies into massive binaries in star forming regions (Sana et al., 2012, 2013; Moe & Stefano, 2013, 2015; Dunstall et al., 2015).
@@ 4.3 HVS candidate selection applied to simulations
Most of the selections described in Section 2.2 can be directly applied to the resulting mock HVS catalogue. We can calculate the implied distance and radial velocity for the mock HVSs in the same way we did for the Gaia catalogue and apply our selections. The same can be done for the photometric, sky coordinate, and instrument-specific selections. An exception is the photo-geometric distance from Bailer-Jones et al. (2021). To apply this selection to the mock HVS catalogue, we resort to the Bailer-Jones (2015) source code (C.A.L. Bailer-Jones, private communication) and calculate the photo-geometric distances for the simulated HVSs. After obtaining the photo-geometric distances for all our mock HVSs in this way, we again apply the same selections to our mock catalogue as to the Gaia catalogue. The only exception is the RUWE, which we do not model for our simulated HVSs. We expect HVSs to be single stars and for single stars fewer than one in a million stars have ${\tt RUWE}>1.4$ (Penoyre et al., 2022).
@@ 5.1 HVSs observable by Gaia
In the left column, we add the contributions from all MFs within our prior range (see Section 4.2). In the middle and right columns we show the distribution for two specific MFs (a top-heavy one, $\kappa=1.7$, and one close to the canonical Salpeter MF, $\kappa=2.3$). The resulting distribution should be seen as a prediction for the overall population distribution of HVSs present in the Gaia catalogue, discovered or undiscovered. We can see that most of the HVSs in the Gaia catalogue will be at heliocentric distances between about 10 and 200 kpc, be on the order of a few solar masses, and be main-sequence stars. Below heliocentric distances of $\sim 5$ kpc, the volume covered by Gaia is too small which makes discovering any HVSs within this region highly unlikely. The vertical feature in the top row at $\sim 8$ kpc is the GC, where, due to projection effects, an over density of HVSs on the sky is present. The cutoff feature towards low-mass/distant stars is caused by the detection limit of Gaia in apparent magnitude. The stars below this boundary are giants, which are visible out to further distances than main-sequence stars. The upper right of the figure is not populated, because extremely massive stars do not live long enough to travel far from the GC. In Appendix Fig. 15, we additionally show the apparent magnitude distribution for a MF index $\kappa=2.3$.
@@
We can see that the completeness for HVSs in Gaia approaches zero for masses below $\sim 1\;{\rm M_{\odot}}$ even when only considering HVSs within 50 kpc from the Sun, which are the ones our observations are mainly sensitive to. For high mass HVSs the completeness converges to nearly one, because those stars do not live long enough to travel far from the GC. because of the effective distance limit imposed by the short life times of massive stars.
@@ 6.4 Distant HVS candidates
We use the posterior on the ejection rate from Fig. 9 to make a prediction on the number of HVSs within the MMT HVS Survey (Brown et al., 2014). We reproduce the same colour selections outlined in section 2.1 of Brown et al. (2012). For simplicity, we consider the footprint of SDSS to be ${\rm Dec}>0$ deg for the Northern Galactic cap, and ${\rm Dec}>-15$ deg for the Southern Galactic cap with a Galactic latitude $|l|>30$ deg (for the sky coverage see Aihara et al., 2011). Brown et al. (2018) only studies stars with a heliocentric radial velocity transformed to the Galactic frame greater than $275$ $\text{km}\,\text{s}^{-1}$, which we follow. In their study, Brown et al. (2018) identify seven stars with a probable origin in the GC. We use the posterior from Fig. 9, in combination with the above selections to calculate how many HVSs there are predicted to be within this footprint. We sample the MF index in the range $[1.7,2.3]$ and multiply each sample by the posterior on the ejection rate for that MF index. We find that within a 68% confidence interval, we expect there to be between $0.3$ and $2.8$ HVSs in the MMT survey footprint. For the 95% confidence interval, we expect there to be between $0.1$ and $5.1$ HVSs respectively. Interestingly, there is thus a $>2\sigma$ tension between our results and the number of GC origin HVSs found in Brown et al. (2018). Apart from a statistical deviation, we have two main hypotheses as to the origin of this tension: the ejection rate of HVSs has not been constant over the flight time of the stars identified in Brown et al. (2018), or (a fraction of) the stars identified in Brown et al. (2018) have a different origin (supported by Generozov & Perets, 2022). To investigate the first hypothesis, we show the distribution of flight times for predicted HVSs in the MMT footprint in Fig. 13 .
@@
Brown et al. (2018) estimated the ejection rate of unbound HVSs in the mass range $2.5\leq M_{\odot}\leq 4$ to be $1.5\times 10^{-6}$ yr-1. Using our observations, we can provide an independent constraint on the ejection rate in this mass range. We do this following the same procedure as for Fig. 10, but now with the above mentioned mass range. Our constraints on this ejection rate are relatively insensitive to the MF varying by a factor of about 3. The $2\sigma$ upper limit peaks at about $1.3\times 10^{-6}$ yr-1 for a MF index $\kappa\sim 2.5$, while the mode of the posterior peaks at about $2.8\times 10^{-7}$ yr-1 (at the same MF index). Our maximum likelihood prediction is thus more than five times lower compared to the findings in Brown et al. (2018).

### v2

# Improved constraints on Galactic Centre ejection of hypervelocity stars based on novel search method

@@ Abstract
Hypervelocity stars (HVSs) are stars which have been ejected from the Galactic Centre (GC) at velocities of up to a few thousand $\text{km}\,\text{s}^{-1}$. They are tracers of the Galactic potential and can be used to infer properties of the GC, such as the initial-mass function and assembly history. HVSs are rare, however, with only about a dozen promising candidates discovered so far. In this work we make use of a novel, highly efficient method to identify new HVS candidates in Gaia. This method uses the nearly radial trajectories of HVSs to infer their distances and velocities based on their position and Gaia proper motion alone. Through comparison of inferred distances with Gaia parallaxes and photometry we identified 600 HVS candidates with G<20 including the previously discovered S5-HVS1, out of which we obtained ground-based follow-up observations for 196 stars. As we found no new HVSs based on their radial velocity, we used detailed HVS ejection simulations to significantly improve previous HVS ejection rate constraints. In particular, the ejection rate of HVSs more massive than 1 $\mathrm{M_{\odot}}$ cannot be higher than $10^{-5}$ yr-1 at $2\sigma$ significance. Additionally, we predict that there are 5–45 unbound HVSs in the complete Gaia catalogue ($1\sigma$ interval), most of which will be main-sequence stars of a few M at heliocentric distances of tens to hundreds of kpc. By comparing our results to literature HVS candidates, we find an indication of either a time-dependent ejection rate of HVSs or a non-GC origin of previously identified HVS candidates.
@@
pubyear: 2024pagerange: Improved constraints on Galactic Centre ejection of hypervelocity stars based on novel search methodD
@@ 2 HVS candidate selection
The main challenge when searching for HVSs is that they are extremely rare; there are for instance no HVSs with velocity greater than 700 km s-1 in the 34M sources in Gaia DR3 with radial velocities, that could be robustly identified (Marchetti et al., 2022). HVS candidates selected from observational catalogues for follow-up observations are therefore typically strongly contaminated by non-HVSs. Here we present our new method to identify HVS candidates originating from the GC. In Section 2.1 we present the general principle we use to find HVSs, and in Section 2.2 we provide the precise selections on the Gaia catalogue for our observational campaign.
@@ 2.1 Radial trajectory
Let $\vec{R}$ be the vector from the GC to the star and $\vec{V}$ the star’s velocity vector relative to the GC. We can write these down in terms of the observables
@@
Figure 1: A diagram of the position ($\vec{R}$) and velocity ($\vec{V}$) of an HVS relative to the GC. This can be expressed in terms of a combination of the star’s observed position ($\hat{n}$) and proper motion ($\vec{\mu}$) and measurements of the Sun’s position relative to the GC ($\vec{R_{0}}$) and velocity ($\vec{V_{0}}$) along with the distance to the star ($D$) and its radial velocity ($V_{r}$).
@@
We convert $D_{\rm I}$ to implied parallax ($\varpi_{\rm I}$) to allow for convenient comparison to measurements by taking the inverse of $D_{\rm I}$. We calculate the uncertainty on $\varpi_{\rm I}$ assuming that the proper motion measurement is the only source of uncertainty. We make the simplifying assumption that the proper motion uncertainties are uncorrelated and since $\varpi_{\rm I}\propto\mu$, we can approximate the uncertainty on $\varpi_{\rm I}$ as
$\sigma_{\varpi_{\rm I}}=\frac{\sqrt{\left[\vec{R}_{0}\cdot\left(\hat{n}\times\hat{\mu}_{\delta}\right)\sigma_{\mu_{\delta}}\right]^{2}+\left[\vec{R_{0}}\cdot\left(\hat{n}\times\hat{\mu}_{\alpha*}\right)\sigma_{\mu_{\alpha}}\right]^{2}}}{\hat{n}\cdot\left(\vec{R}_{0}\times\hat{V}_{0}\right)},$ (7)
with $\hat{\mu}{{}_{\delta}}$ and $\hat{\mu}_{\alpha}$ the basis vectors for the proper motion in declination (Dec) and right ascension (RA) respectively, and $\sigma_{\mu_{\delta}}$ and $\sigma_{\mu_{\alpha}}$ the uncertainties on the proper motions in Dec and RA respectively. For a more accurate uncertainty analysis one would sample over the uncertainty on all observables.
@@ 2.2.1 Astrometric and kinematic selections
We start by using the astrometric and kinematic properties of stars to select potential HVSs. HVSs need to be very fast to prevent significant deviation from their otherwise radial trajectories by the torque on the orbit generated by the non-sphericity of the Galactic potential. For this reason, we only consider stars with total implied velocity, $V_{\text{I}}$, in the range $[800,3500]$ $\text{km}\,\text{s}^{-1}$, independent on Galactocentric distance. The lower limit ensures that (excluding the inner GC) all stars would be unbound. The upper limit includes most of the extremely fast predicted HVSs (Rossi et al., 2014), while limiting contamination by non-HVSs (which can have extremely high implied velocities). Secondly, we compare the implied parallax to the parallax measured by Gaia. We only select sources with implied parallax consistent with measured parallax within $2\sigma$, considering both the uncertainty on the measured and the implied parallaxes. We do this comparison to the parallax and not the distance since the parallax has Gaussian uncertainties.
@@ 2.2.3 Sky coordinate selections
We know that the distribution of HVSs on the sky should be anisotropic, even for an isotropic ejection mechanism in the GC due to on-sky projection. Moreover, we expect that the contamination of our HVS candidate sample is highest near the GC and anti-centre, where our selections are relatively ineffective. In addition, in the Galactic plane we expect the contamination to be higher than at high Galactic latitudes due to the high relative number of field stars to expected HVSs. In order to limit contamination by field stars we define the expected number of HVSs to number of candidates for different directions in the sky. For this purpose we use HEALPix [^2] http://healpix.sf.net/ ($nside=3$; Górski et al., 2005; Zonca et al., 2019) to divide the sky into equal area pixels. We determine the expected density of HVSs across the sky using the simulations described in Section 4. We only observe sources in HEALPix pixels where the ratio of the density of expected HVSs to candidates (normalised by the number of candidates) is larger than five [^3] The HEALPix pixels we use are 0-19, 22-31, 35-42, 46-55, 58-62, 65, 66, 69-73, 78, 82-85, 93-96, 101, 102, 103 and 107 in the ring scheme based on RA and Dec..
@@
The selection steps up to this point leave us with only 600 HVS candidates, the previously identified S5-HVS1 being one of the top candidates (Koposov et al., 2020). We provide an overview of the selections in Table 1.
@@ 3.4 Observed HVS candidates
Figure 3: Sky distribution of our HVS candidate sample in Galactic coordinates. Black points indicate the stars with follow-up observations and the orange points the ones with either SDSS or LAMOST radial velocity measurements. For the grey stars we do not have radial velocity measurements.
@@
In Fig. 4 we plot the implied radial velocity against the measured radial velocity.
@@ 4.1 Hills mechanism
where we neglected a multiplicative constant factor of order unity which depends on the geometry of the encounter (see, e.g., Fig.10 in Sari et al., 2009). In equation 8, $m_{\text{c}}$ is the mass of the captured companion, $M_{\text{MBH}}$ in the case of the Galaxy is the mass of Sgr A ($4\times 10^{6}$ M, Eisenhauer et al., 2005; Ghez et al., 2008; Schödel et al., 2009; Gillessen et al., 2009; Boehle et al., 2016; Gillessen et al., 2017; Do et al., 2019; GRAVITY Collaboration et al., 2019), and $M$ the total mass of the progenitor binary (Sari et al., 2009; Kobayashi et al., 2012; Rossi et al., 2014). Most of the stars ejected through this mechanism are still bound to the Galaxy, but the ejections can occur at velocities in excess of 1000 $\text{km}\,\text{s}^{-1}$, putting stars on orbits unbound to the Galaxy.
@@ 4.2 Simulations
We start by creating an HVS progenitor population of binaries. These binaries are characterised by the primary zero-age main sequence mass ($m_{\text{p}}$), mass ratio ($q$), and the orbital semi-major axis ($a$) at the moment of disruption. The mass of the primary is drawn from a mass function (MF) which follows a power law distribution of the form $f(m_{\text{p}})\propto m_{\text{p}}^{-\kappa}$ in the range $[0.1,100]$ M. The secondary is then assigned a mass according to a mass ratio distribution of the form $f(q)\propto q^{\gamma}$, with $0.1\leq q\leq 1$. Lastly, the orbital semi-major axis are assigned assuming a binary orbital period distribution of the form $f(\log P)\propto(\log P)^{\pi}$. Different values for the three parameters $\kappa$, $\gamma$, and $\pi$ are explored with uniform priors to account for uncertainty in the HVS progenitor population. For $\kappa$ we use the interval $[0.3,3.3]$ to allow for both very top-heavy and bottom-heavy MFs. For $\gamma$ we use the interval $[-2,2]$ and for $\pi$ the interval $[0,2]$. We allow for the wide range in MF index because of the uncertainty on the IMF near the GC, alluded to in the introduction. IMF indexes of 0.45, 1.7, and a canonical 2.3 have all been claimed near Sgr A (Bartko et al., 2010; Lu et al., 2013; Löckmann et al., 2010, respectively). The parameter range on $\pi$ allows for a log-uniform period distribution (also known as Öpik’s law), or distributions with increasing fractions of wide binaries. The parameter range on $q$ encapsulates the results from different studies into massive binaries in star forming regions (Sana et al., 2012, 2013; Moe & Stefano, 2013, 2015; Dunstall et al., 2015).
@@ 4.3 HVS candidate selection applied to simulations
Most of the selections described in Section 2.2 can be directly applied to the resulting mock HVS catalogue. We can calculate the implied distance and radial velocity for the mock HVSs in the same way we did for the Gaia catalogue and apply our selections. The same can be done for the photometric, sky coordinate, and instrument-specific selections. An exception is the photo-geometric distance from Bailer-Jones et al. (2021). To apply this selection to the mock HVS catalogue, we resort to the Bailer-Jones (2015) source code (C.A.L. Bailer-Jones, private communication) and calculate the photo-geometric distances for the simulated HVSs. After obtaining the photo-geometric distances for all our mock HVSs in this way, we again apply the same selections to our mock catalogue as to the Gaia catalogue. The only exception is the RUWE, which we do not model for our simulated HVSs. We expect HVSs to be single stars and for single stars fewer than one in a million have ${\tt RUWE}>1.4$ (Penoyre et al., 2022).
@@ 5.1 HVSs observable by Gaia
In the left column, we add the contributions from all MFs within our prior range (see Section 4.2). In the middle and right columns we show the distribution for two specific MFs (a top-heavy one, $\kappa=1.7$, and one close to the canonical Salpeter MF, $\kappa=2.3$). The resulting distribution should be seen as a prediction for the overall population distribution of HVSs present in the Gaia catalogue, discovered or undiscovered. We can see that most of the HVSs in the Gaia catalogue will be at heliocentric distances between about 10 and 200 kpc, be on the order of a few solar masses, and be main-sequence stars. Below heliocentric distances of $\sim 5$ kpc, the volume covered by Gaia is too small which makes discovering any HVSs within this region highly unlikely. The vertical feature in the top row at $\sim 8$ kpc is the GC, where, due to projection effects, an over density of HVSs on the sky is present. The cutoff feature towards low-mass/distant stars is caused by the detection limit of Gaia in apparent magnitude. The stars below this boundary are giants, which are visible out to further distances than main-sequence stars. The upper right of the figure is not populated, because extremely massive stars do not live long enough to travel far from the GC. In Appendix Fig. 15, we additionally show the apparent magnitude distribution for am MF index $\kappa=2.3$.
@@
We can see that the completeness for HVSs in Gaia approaches zero for masses below $\sim 1\;{\rm M_{\odot}}$ even when only considering HVSs within 50 kpc from the Sun, which are the ones our observations are mainly sensitive to. For high mass HVSs (M $\gtrsim 5$ M) the completeness converges to nearly one, because those stars do not live long enough to travel far from the GC.
@@ 5.2 HVS parameter constraints
Overall, our upper limit ejection rate constraints are consistent with the process where stellar binaries’ centre of mass trajectories have undergone gravitational scatterings off other (single) stars or giant molecular clouds: once every $10^{4}-10^{6}$ yr a binary star is sent on a highly eccentric orbit such that its pericenter becomes equal or smaller than the tidal separation radius (a.k.a. “loss cone orbit”) and an HVS is ejected (Yu & Tremaine, 2003; Perets et al., 2007). For a recent review on rates see Stone et al. (2020, especially section 3.6.3). Robust rate predictions to compare with our observational constraints should include realistic descriptions of the phase space and properties of binary stars, as well as those of single stars and giant molecular clouds out to $\sim 100$ pc from SgrA* in our GC. Such calculations and comparisons are beyond the scope of this paper and will be presented in a follow up investigation.
@@
In the hypothetical case we observed two HVSs instead of one, the effect on our constraints is mainly that the posterior (Figs. 9 and 10) becomes more narrow (the $2\sigma$ confidence interval by $\sim 0.5$ dex). In addition, the $2\sigma$ upper limit constraint would increase by about $\sim 0.1$ dex.
@@ 6.4 Distant HVS candidates
We use the posterior on the ejection rate from Fig. 9 to make a prediction on the number of HVSs within the MMT HVS Survey (Brown et al., 2014). We reproduce the same colour selections outlined in section 2.1 of Brown et al. (2012). For simplicity, we consider the footprint of SDSS to be ${\rm Dec}>0$ deg for the Northern Galactic cap, and ${\rm Dec}>-15$ deg for the Southern Galactic cap with a Galactic latitude $|l|>30$ deg (for the sky coverage see Aihara et al., 2011). Brown et al. (2018) only studies stars with a heliocentric radial velocity transformed to the Galactic frame greater than $275$ $\text{km}\,\text{s}^{-1}$, which we follow. In their study, Brown et al. (2018) identify seven stars with a probable origin in the GC. We use the posterior from Fig. 9, in combination with the above selections to calculate how many HVSs there are predicted to be within this footprint. We sample the MF index in the range $[1.7,2.3]$ and multiply each sample by the posterior on the ejection rate for that MF index. We find that within a 68% confidence interval, we expect there to be between $0.3$ and $2.8$ HVSs in the MMT survey footprint. For the 95% confidence interval, we expect there to be between $0.1$ and $5.1$ HVSs respectively. It is thus likely that a number of the stars identified in Brown et al. (2018) are genuine HVSs. Interestingly however, there is thus a $>2\sigma$ tension between our results and the number of GC origin HVSs found in Brown et al. (2018). Apart from a statistical deviation, we have two main hypotheses as to the origin of this tension: the ejection rate of HVSs has not been constant over the flight time of the stars identified in Brown et al. (2018), or (a fraction of) the stars identified in Brown et al. (2018) have a different origin (supported by Generozov & Perets, 2022). To investigate the first hypothesis, we show the distribution of flight times for predicted HVSs in the MMT footprint in Fig. 13 .
@@
Brown et al. (2018) estimated the ejection rate of unbound HVSs in the mass range $2.5\leq M_{\odot}\leq 4$ to be $1.5\times 10^{-6}$ yr-1. Using our observations, we can provide an independent constraint on the ejection rate in this mass range. We do this following the same procedure as for Fig. 10, but now with the above mentioned mass range. Our constraints on this ejection rate are relatively insensitive to the MF varying by a factor of about 3. The $2\sigma$ upper limit peaks at about $1.3\times 10^{-6}$ yr-1 for an MF index $\kappa\sim 2.5$, while the mode of the posterior peaks at about $2.8\times 10^{-7}$ yr-1 (at the same MF index). Our maximum likelihood prediction is thus more than five times lower compared to the findings in Brown et al. (2018).